Two cars leave from the same place at the same time, traveling in opposite directions. After 2 hours, the cars are 308 km apart. What are their speeds if their speeds are in the ratio 4:3?
Answer by mananth(16949)

The ratio of their speeds = 4/3
multiply by a common factor x
their speeds be 3x & 4x
opposite direction so add speeds= 7x
T= d/r
308/7x=2
x=308/14
x=22
..
one's apeed = 66 km/h
the other's speed =88km/h